OpenStudy (anonymous):

Ammonia, NH3, is a typical ingredient in household cleaners. It is produced through a combination reaction involving N2(g) and H2(g). If 24.0 mol of H2(g) react with excess N2(g), how many moles of ammonia are produced? Show your work. Please help im really confused

OpenStudy (aaronq):

for problems involving stoichiometry, you have to convert whatever information they give you to moles (signified by the symbol "n") Build a ratio using the stoichiometric coefficients, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{red}{b}B\) \(\rightleftharpoons\) \( \color{red}{c}C\) where upper case are the species, and lower case are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{red}{b}}=\dfrac{n_C}{\color{red}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

okay, I understand the whole concept of what you are saying.. however im still a bit confused on how i would set up the, Ng & Hg. you explained yourself very well, thank you. but can you help me on where i need to set these up and solving the 24.0 please, thank you again. @aaronq

OpenStudy (aaronq):

Okay, so they tell you there is an excess of \(N_2\), meaning the limiting reactant (that which determines the extent of the reaction) is \(H_2\). so, i forgot to mention that you have to write an equation for the reaction first, and balance it. Do you think you can do that?

OpenStudy (anonymous):

I believe so,@aaronq will you tell me if i make a mistake or not?

OpenStudy (aaronq):

Yep, i will.

OpenStudy (anonymous):

Okay so what i'm understanding is i'm using 24.0 moles of H2. And Nh3 is producing the combination of N2 and H2 so therefore making the equation 3H2 +N2 = 2N H3 okay so with using 24 moles from H2 which is now H3 24/3 =8 and using 8 being multiplied into 2 because 2N now gives you 24/3=8 8 x 2 =16 okay im kinda confused now lol.. but is 16 the amount of ammonia produced? @aaronq

OpenStudy (anonymous):

Please tell me this is right, i watched youtube videos to help me understand a little better.. so im hoping this is the right result

OpenStudy (aaronq):

when you use the ratios, you take the molecule as a whole. By that, i mean you don't split \(NH_3\) into N and \(H_3\). \(\color{red}3H_2 +N_2 \rightleftharpoons \color{blue}2N H_3\) \(\dfrac{n_{H_2}}{\color{red}3}=\dfrac{n_{NH_3}}{\color{blue}2}\) from here, plug in your information: \(\dfrac{24\;mol}{\color{red}3}=\dfrac{n_{NH_3}}{\color{blue}2}\) Then you would solve this by isolating \(n_{NH_3}\) algebraically.

OpenStudy (jtfan2000):

8 mol NH3

OpenStudy (aaronq):

@JTfan2000 nah 16 mol

OpenStudy (jtfan2000):

@aaronq - yes, 8 mol - Ammonia, NH3, is a typical ingredient in household cleaners. It is produced through a combination reaction involving N2(g) and H2(g). If 12.0 mol of H2(g) react with excess N2(g), how many moles of ammonia are produced? Show your work.

OpenStudy (aaronq):

@JTfan2000 Except the original question read "If \(\sf 24.0\) mol of H2(g) react with excess N2(g), how many moles of ammonia are produced?"